A survey conducted in Mega Manila with a sample of 500 gasoline stations found that the price charged per liter of gasoline is normally distributed with a mean of P55 and a standard deviation of P3. Using the empirical rule, determine the number of stations that charge between P50 and P54.

1
Expert's answer
2021-11-09T11:25:31-0500

"n=500 \\

\mu=55 \\

\sigma= 3 \\

P(50<X<54) = P(X<54) -P(X<50) \\

=P(Z< \frac{54-55}{3}) -P(Z< \frac{50-55}{3}) \\

= P(Z< -0.333) -P(Z< -1.666) \\

= 0.3695 -0.0478 \\

= 0.3217 \\

N = 500 \times 0.3217 = 160.85 ≈ 161"

The number of stations that charge between P50 and P54 is 161.